### v1

# Joint Learning for Scattered Point Cloud Understanding with Hierarchical Self-Distillation

Abstract
Numerous point-cloud understanding techniques focus on whole entities and have succeeded in obtaining satisfactory results and limited sparsity tolerance. However, these methods are generally sensitive to incomplete point clouds that are scanned with flaws or large gaps. To address this issue, in this paper, we propose an end-to-end architecture that compensates for and identifies partial point clouds on the fly. First, we propose a cascaded solution that integrates both the upstream and downstream networks simultaneously, allowing the task-oriented downstream to identify the points generated by the completion-oriented upstream. These two streams complement each other, resulting in improved performance for both completion and downstream-dependent tasks. Second, to explicitly understand the predicted points’ pattern, we introduce hierarchical self-distillation (HSD), which can be applied to arbitrary hierarchy-based point cloud methods. HSD ensures that the deepest classifier with a larger perceptual field and longer code length provides additional regularization to intermediate ones rather than simply aggregating the multi-scale features, and therefore maximizing the mutual information between a teacher and students. We show the advantage of the self-distillation process in the hyperspaces based on the information bottleneck principle. On the classification task, our proposed method performs competitively on the synthetic dataset and achieves superior results on the challenging real-world benchmark when compared to the state-of-the-art models. Additional experiments also demonstrate the superior performance and generality of our framework on the part segmentation task.
Index Terms:
Joint Learning, hierarchical self-distillation, scattered point cloud understanding, information bottleneck.I Introduction
Point cloud completion and understanding are essential in quite a number of 3D computer vision tasks, including autonomous driving, augmented reality, industrial manufacturing, and robotic maneuver. Especially, in interactive scenarios such as robotic grasp, an ill-scanned object or scene leads to difficulty on accurate interactions for two reasons. First, planes and objects need to be reconstructed from sensor-produced point clouds for collision detection and grasp pose selection. Second, the interaction manner and force are dependant upon the cognition of objects that is subject to the irregularity and sparseness of point clouds. In recent years, numerous studies [1, 2, 3] were proposed to leverage point cloud directly for understanding. While the globally- and locally- extracted features increase the tolerance of irregularity and sparseness, they are hindered from processing point clouds with a large amount of missing parts or overly sparse surfaces due to fast scanning, occlusion, view point, or sensor resolution. As a result, [4, 5] were proposed specifically concentrating on sparse point clouds to tackle this problem. However, these methods utilizing either sophisticated voxelization procedure or multi-view photometric information do not generalize well due to limited memory or the need of additional supervisory signal.
An intuitive plan is to first compensate the incomplete parts, and then to sequentially apply a recognition algorithm on the completed point clouds. Nevertheless, models trained separately may suffer from lower robustness caused by the inconsistency in point distributions. Without explicitly learning the overall shape, as illustrated in Fig. 1(a), the network may allocate excessive attention to irrelevant points or outliers caused by irregular local shapes, leading to degraded performance. The masked autoencoder (MAE) was initially developed to pre-train encoders by learning the correspondence between masked sections and original input data such as texts, images, or point clouds. This facilitated fine-tuning processes for natural language processing (NLP) [6] and computer vision [7, 8, 9] tasks. Albeit MAE is primarily regarded as a pre-training model for improving classification performance, its reconstruction capacity is advantageous in scenarios where observations are incomplete. Additionally, hierarchy-based networks are designed to aggregate sparse features learned from multiple scales for better coverage of local regions. However, as shown in in Fig. 1(b) (left), the aggregation operation may result in the aggregated representation being unaware of the mutual information across scales. As a solution, Deeply-Supervised Nets (DSN) [10] provides separate guidance to each scale as a strong regularization technique, but is still unaware of the correlation among those scales.
In order to tackle the challenges inherent in scattered point cloud classification, we propose an approach that jointly learns the completeness and shape category with hierarchical self-distillation (HSD). It consists of two cascaded streams: upstream and downstream. In the upstream, we generate a small number of portions of a point cloud based on randomly selected centroids that serve as the unmasked parts. Thereafter, a MAE is leveraged to learn the asymmetric mapping from the unmasked (complete) and masked (incomplete) point clouds. In the downstream, we can use any hierarchy-based point cloud classifier, such as PointNet++ [2], CurveNet [11], or PointMLP [12]. To further regularize the training process, our HSD framework aims to maximize the mutual information from multiple coverage scales and transfers the most distinct features in the deepest hyperspace to other scales, as shown in Fig. 1(b) (right). Specifically, the last classifier in the series from different scales is designated as the teacher to supervise the previous ones, thus transferring the important knowledge to different hyperspaces. We show that DSN can be viewed as a special case of our HSD. Theoretically, we leverage the information bottleneck principle to explain our model and demonstrate its generalization performance in experiments.
Our method places emphasis on joint learning, where both MAE and HSD are leveraged to harmonize distributions, thus enhancing the efficacy of both tasks, i.e., reconstruction and understanding. Our primary contributions are summarized as follows:
An end-to-end partial point cloud processing network is proposed to jointly learn the geometric features from irregular shapes, which dedicates to leverage the reconstruction with low-level shapes to help understanding, and in turn, to use understanding of high-level details to help the reconstruction.
A hierarchical self-distillation (HSD) framework is proposed to maximize the mutual information captured in scale-aware hyperspaces. Its plug-and-play property is highly flexible and therefore can be integrated to any hierarchy-based point cloud baselines.
Our evaluation results on scattered point clouds demonstrate that our method performs competitively on ModelNet40, while also attains state-of-the-art (SOTA) performance on the real-world dataset ScanObjectNN for classification, as well as ShpeNetPart for part segmentation.
II Related Work
Masked Autoencoder. Initially, the masked autoencoder (MAE) was introduced as a self-supervised solution to handle the lack of accessible labels in data for NLP [6]. Although the information density in computer vision is different from that of NLP, researchers have shown that MAE works effectively in computer vision by masking out certain portions of pixels in the asymmetric architecture to effectively learn largely missing patches in the pixel space [7, 8, 9, 13]. Following the image works, PointMAE [14] was the first MAE method for point clouds. Based on PointMAE, more 3D related MAEs [15, 16] have been proposed.
These methods, however, treat MAE as a pre-training model, requiring a fine-tuning stage for new data prediction. Our aim is to tackle incomplete object classification in an end-to-end paradigm, requiring only the process of self-supervised reconstruction in MAE. The completion networks [17, 18, 19, 20, 21, 22, 23] are well-suited for this purpose, as the synthetically pre-processed incomplete data can be viewed as naturally masked point clouds. SnowflakeNet [21], operating on point splitting, achieves satisfactory reconstruction results with a relatively simple structure. In this study, we adopt the simplified SnowflakeNet as a masked autoencoder for reconstruction by masking out random areas on the input point set.
Conventional Knowledge Distillation. Aiming at improving the performance of deep learning models, a common way is to average all predictions on the same data. This cumbersome scheme expedites the invention of conventional knowledge distillation that transfers the knowledge among different models [24]. Generally, teacher networks with greater model complexity learn knowledge and guide student networks with more compactness to learn from better representations for enhanced generalization. As a result, student networks are more suitable for deployment on systems with less computational resource [25, 26], e.g., mobile phone and circuit board. Other investigations into knowledge distillation, with an emphasis on early stopping of teacher’s training, have asserted that fully converged teachers do not typically promote student learning [27, 28]. In this study, we trivially bypass the need for optimal teacher selection through end-to-end self-distillation.
Self-Distillation. Born-again networks [29] revisited knowledge distillation with a different objective by disentangling the training procedure from model compression. In a progressive fashion, the student models that are identical to the teacher model are trained consecutively from the supervisions of previous generations, forming an ensemble at the end that even outperforms the teacher model at the first generation.
Recently, a new concept called self-distillation has emerged following the development of born-again networks. This can be divided into two categories based on whether knowledge is distilled within [30, 31, 32] or out of [33, 34] the same iteration in the training stage, for tasks such as label smoothing or weakly supervised learning perturbation. Our proposed method belongs to the former one, in which we train each classification model simultaneously regardless of the roles. Inspired by those born-again networks, we tend to transfer the knowledge from the deepest level of the hierarchy to shallower models through self-distillation.
Point Cloud Understanding. Studies on point cloud have demonstrated outstanding performances on the publicly recognized benchmarks. Pioneered by PointNet [1], hierarchical scaling networks were proposed to enhance the global and local awareness, including PointNet++ [2], DGCNN [3], GBNet [35], CurveNet [11], PointMLP [12], Point-PN [36] and PointConT [37]. PointNet++ and DGCNN formulate the hierarchical structures by learning the point relation from multi-scale grouping (MSG) or k nearest neighbors (kNN) in the local region. CurveNet walks along the sequence of connected points, forming hypothetical curves that augment pointwise features. PointMLP rethinks PointNet++ with simple yet effective residual connections, holding the state of the art. In this paper, we meticulously select several well-established hierarchy-based networks (i.e., PointNet++, CurveNet, PointMLP) in our joint learning framework for the downstream task.
III Method
Problem Formulation. Given the complete point set and its incomplete counterpart , we seek for a solution to reconstruct from and simultaneously classify/segment the reconstructed , where , , and denote the number of each point set in the Euclidean space, respectively. The architecture of our method is illustrated in Fig. 2. Our end-to-end architecture is composed of upstream and downstream sub-networks, denoted as and .
III-A Preliminary
Information Bottleneck. Let the random variable represent the input data, represent the output/target labels, the information bottleneck (IB) principle [38] aims to find a compressed representation of , denoted by , that minimizes the mutual information and maximizes the relevant information . This objective can be formulated as:
where is a Lagrange multiplier that constrains the expected distortion.
Strong Regularization. To solve the gradient vanishing problem, Deeply-Supervised Nets (DSN) [10] simultaneously minimizes loss functions across hidden layers [39]. DSN functions as a strong regularization and it smooths the gradient in the back propagation. From the perspective of IB principle, the optimization of DSN can be formulated as:
where controls the emphasis on preserving relevant information. In a cascaded manner, each layer performs analogous to an ensemble of previous layers (companions). Obviously, the second term can be regarded as the regularization over a conventional single objective function. Intuitively, DSN treats both the output layer and companions equally.
III-B Upstream Network
In real scenes, it is common that the scanned point clouds are incomplete or imperfect due to sensor resolution, obstacles, or vibration. To alleviate this phenomenon, we adopt a simplified version of SnowflakeNet to be our upstream network to learn the mapping between and , which is optimized by a reconstruction loss:
where denotes the squared Euclidean 2-norm of Chamfer distance (CD).
III-C Downstream Network
Generally, the downstream network is dependent upon varying tasks. Targeting on point cloud classification problem, we leverage three elite versions of canonical hierarchy-based methods, i.e., PointNet++, CurveNet, and PointMLP, as our downstream networks, which have demonstrated satisfactory local awareness. The objective in this task obeys Eq. 2, which is cross-entropy loss, denoted as . Without losing generality, most point cloud processing methods have aggregation operations (e.g., max pooling, summation, and global set abstraction) at the end of feature extraction, because of the orderless nature of point clouds. However, such operations potentially eliminate the distinguishable features, which leads to degraded performances. To effectively learn the knowledge across different scales, we propose to use hierarchical self-distillation as a novel feature fusion framework, which will be described in the next section.
III-D Stronger Regularization
Hierarchy-based networks are capable of producing multi-level features that are generated from different scales, e.g., the integrated features from MSG of PointNet++ in Fig. 1(b). However, MSG-like schemes have limitations, as they do not capture the dissociated information among features. Hence, we propose a solution that involves building multiple classification heads for hierarchical feature extraction and utilizing self-distillation to learn comprehensive knowledge among all levels. We dub this scheme as PointHSD, a general framework that can be employed by arbitrary hierarchy-based point cloud methods. A universal structure for hierarchical feature extraction (HFE) that includes the elite versions of classical point cloud classification networks (i.e., PointNet++, CurveNet, and PointMLP) is illustrated in Fig. 3. Here, set abstraction, curve convolution, and residual MLPs serve as local feature extractors that already embed local aggregation layers inside. For further details, please refer to the supplementary material.
The Hierarchical Feature Extraction (HFE) module, as shown in Fig. 2, is responsible for learning the high-level local details from the low-level reconstructed point set . This is achieved by using different branches, each of which represents a distinct range of surroundings. Specifically, let denote this procedure, yielding intermediate and deepest latent representations and . We want to have a larger perceptual field, such that the difference and between and intermediate can be learned by self-distillation. The regularization term in Eq. 2 is therefore reformulated as:
By maximizing the mutual information , we can equivalently quantify the difference (denoted as ) by minimizing the Kullback-Leibler (KL) divergence. The third term in Eq. 4 further regularizes the optimization by smoothing the hard predicted labels of students (Fig. 1(b)). As such, our HSD framework treats the output layer and companions distinctively. Apparently, deep supervision [10] can be seen as a special case of Eq. 4 if .
Lemma If a teacher has a larger local perceptual field and code length than students, then .
Proof Assume for any , holds, then:
Since , we have . Along with the cross-entropy term pertaining to students, the conditional mutual information (equivalently quantified by ) imparts smoothness to the probabilistic distribution through the convolution of information derived from the teacher. This mechanism contributes to stronger regularization of the students, thereby ultimately mitigating the overconfidence of the ensemble model.
III-E Joint Learning
Recall the intuitive example in Fig. 1(a), as the resolution or completeness of input point clouds declines, the accuracy of classification methods tend to decrease substantially. This is due to the fact that incomplete shapes can lead to misleading local features.
In order to overcome this issue, we introduce a simple yet effective function to jointly learn the completeness and category of objects, denoted as:
where means the cascade of two streams, and follows the Markov chain with the assumption that . Therefore, we designate the deepest layer as the teacher since it contains most comprehensive knowledge to smooth the hard labels in the companions. Mutual information is implicitly compressed by the network. Finally, combining Eqs. 2, 3, and 4, our joint learning objective is optimized by
where , , are the weights, which are empirically set to 1, 0.001, and 0.8, respectively.
It is important to note that our joint learning framework shares similarities with the cascaded fusion network, CasFusionNet [40], designed for scene completion and semantic segmentation. The fundamental distinction lies in that CasFusionNet focuses on dense feature fusion within the hyperspace, amalgamating both geometric and semantic features, while our method extends its influence through the integration of both low-level and high-level features to achieve joint promotion. Our work also bears similarities with GBNet [35] that leverages both low-level and high-level geometric cues. However, the substantial departure lies in that GBNet still simply aggregates local contexts, whereas our HSD perceives and distills dissociated information through features from different scales, as illustrated in Fig. 1(b). Furthermore, our joint learning design imparts a unique capability to our pipeline, making it particularly effective in handling scattered point clouds.
IV Experiment
Dataset. Our data is based on three commonly used benchmarks, namely ModelNet40 [41], ScanObjectNN [42], and ShapeNetPart [43]. ModelNet40 is a synthetic mesh dataset consisting of 9,843 training and 2,468 testing samples, categorized into 40 entities. On the other hand, ScanObjectNN is a challenging real-world dataset with 15,000 objects classified into 15 categories, comprising 2,902 unique object instances. To generate scattered point clouds, instead of randomly masking out points, we sparsely sample a small number of points to be patch centroids using the farthest point sampling (FPS) algorithm, and subsequently attach the surrounding points using the kNN algorithm. Especially, we have three data types as input for ModelNet40: with the size of , , and (the 8 neighbors are sampled based on a dense resolution of the ground truth (GT), making it much sparser than the previous two types), respectively. For ScanObjectNN and ShapeNetPart, we randomly sample points as input.
| Dataset | Resolution | Sparsity |
| ModelNet40 | Low | |
| Medium | ||
| High | ||
| ScanObjectNN | Medium | |
| Medium | ||
| ShapeNetPart | Medium |
Data Preparation. We sample the scattered point clouds based on specific number of seeds and neighbors based upon FPS and kNN. As such, we list the sparsity types for all datasets we have used in Table I. Moreover, in Fig 4, we visualize the Sparsity for different types of inputs for experiments on ModelNet40. Specifically, the input with points has the most scattered distribution as the neighbors are sampled from a dense version, i.e., 10,000 points. The data pre-processing code for all datasets is included in the source code.
Implementation Detail. Our networks are trained on a server deployed with 10 Nvidia RTX 2080Ti GPUs, with a batch size of 260 for classification and 160 for segmentation during training and testing for all models. We do not apply the voting strategy throughout experiments. Except otherwise noted, we omit the upstream name (i.e., SnowflakeNet) and directly use the downstream model names to denote two-stage methods, and use for our joint learning framework.
Evaluation Metrics. Following the standard in the literature, we report the overall accuracy (OA), mean-class accuracy (mAcc), and Chamfer distance (CD) () for classification, and mean-instance and mean-class intersection over union (mIOU/cIOU) for segmentation.
IV-A Classification with Self-Distillation
| Model | OA(%) | mAcc(%) |
| PointNet++ [2] | 91.8 | 89.2 |
| CurveNet [11] | 92.5 | 89.4 |
| PointMLP [12] | 92.7 | 89.7 |
| PointNet++-HSD (Ours) | 92.6 | 89.5 |
| CurveNet-HSD (Ours) | 92.6 | 90.0 |
| PointMLP-HSD (Ours) | 93.2 | 90.3 |
In this study, we investigate the performance of PointNet++, CurveNet, and PointMLP as classification-only models. To demonstrate the difference between these methods and our proposed HSD framework, we evaluate their accuracy on all downstream networks with full input, as presented in Table II, indicating that HSD is capable of refining the classification results. Since the input point cloud is complete, we only take one view of it for this task. To ensure a fair comparison, we re-train all baselines on the same machine as the proposed models. The data is exactly same as the one used in other canonical models: we take 1,024 points as input and the training/testing splitting follows the standard criterion in the community. The best testing results of OA and mAcc are reported. In this task, we utilize the cosine annealing scheduler to establish the learning rate, as same as PointMLP. The minimum learning rate is 0.005.
IV-B Qualitative Visualization
Figs. 5, 6, and 7 visually demonstrate the improvements in the quality of the reconstructed point clouds for ModelNet40, ScanObjectNN, and ShapeNetPart, respectively. To enhance visualization, we employ the PB_T50_RS without background for ScanObjectNN in Fig. 6. Notably, some input objects are unrecognizable even for human, highlighting the necessity of the upstream stage for better understanding of 3D objects. Particularly, in Fig. 5, the chair’s leg can be essential to the final classification, and our HSD methods successfully reconstruct this part. In Fig 7, the quality of reconstructed point clouds is even better than the insufficiently sampled GT. Since the final objective is object recognition rather than completion, the reconstructed point clouds are restricted to have 1,024 points to ensure downstream efficiency, which generally recovers only the overall shape of an incomplete point cloud.
| Model | OA(%) | mAcc(%) | CD | OA(%) | mAcc(%) | CD | OA(%) | mAcc(%) | CD |
| PointNet++ [2] | 89.0 | 88.7 | 23.04 | 87.0 | 86.6 | 27.33 | 70.3 | 70.1 | 40.78 |
| DGCNN [3] | 79.9 | 72.4 | 77.6 | 71.8 | 65.3 | 57.6 | |||
| PCT [44] | 88.9 | 85.0 | 86.6 | 81.6 | 71.7 | 64.8 | |||
| CurveNet [11] | 84.6 | 79.1 | 82.1 | 76.0 | 69.0 | 60.0 | |||
| PointMLP [12] | 85.7 | 78.5 | 81.7 | 72.9 | 66.9 | 59.6 | |||
| Our (PointNet++) | 89.9 | 86.2 | 23.22 | 89.1 | 85.0 | 27.41 | 82.7 | 77.4 | 38.10 |
| Our (CurveNet) | 89.0 | 86.0 | 23.03 | 88.8 | 85.8 | 27.79 | 80.4 | 76.1 | 38.71 |
| Our (PointMLP) | 89.9 | 87.1 | 23.15 | 89.4 | 85.6 | 27.49 | 79.7 | 75.7 | 39.67 |
| Our (PointNet++-HSD) | 90.8 | 87.7 | 23.10 | 89.9 | 86.6 | 27.79 | 82.9 | 76.9 | 39.10 |
| Our (CurveNet-HSD) | 89.1 | 85.5 | 23.16 | 88.8 | 85.3 | 28.02 | 81.3 | 75.5 | 39.17 |
| Our (PointMLP-HSD) | 90.3 | 87.0 | 23.27 | 88.9 | 85.0 | 28.36 | 81.8 | 75.8 | 39.42 |
IV-C Incomplete Point Cloud Classification
The point sets collected by real sensors (e.g., Lidar) are only distributed on the object’s surface and only near the side of Lidar, resulting in misalignment of classification and uncertainty of boundary detection. The upstream network tackles this challenge by reconstructing the incomplete point cloud into a complete shape. In this task, all of our models utilize incomplete point clouds as input. All two-stage models are built with full architectures, while our downstreams are elite versions. The size of the target point cloud is 1,024 for fair comparisons. During training, there are 8 views for each incomplete input. In this task, we gradually decay the learning rate for every 50 epochs with a factor of 0.8. The initial learning rate is 0.001. In following experiments, we denote as our joint learning methods.
Results on ModelNet40. The results in Table III reflect that the proposed HSD models outperform the baselines in terms of OA for incomplete object classification. We also conduct completion and classification tasks separately, rather than in an end-to-end scheme. The results generated by SnowflakeNet are directly fed into the fully trained downstream classification networks. For PointNet++, due to unstable predicted results, we run the evaluation for 4 times and report the average for both OA and mAcc. Our HSD models achieve the best OA with competitive CD compared to the two-stage methods. It can be also observed that our proposed methods significantly outperform the two-stage methods on the sparsely masked data. We infer that our joint learning models enhance the information gain especially on sparsely distributed data, mimicking the inference procedure in the real scenario. It can be observed that HSD works less effective on CurveNet. Empirically, despite CurveNet has a hierarchical nature, it walks among the points, forming the curves to represent local features. Unlike query ball and kNN in conventional hierarchy-based methods, curves share more discriminative features in the same region, making different heads share less mutual information.
Results on ScanObjectNN. In this task, we follow the standard in the community and use the hardest variant of ScanObjectNN: PB_T50_RS with background. In Table IV, although our baseline models (i.e., without HSD) have already achieved competitive performances, our HSD models achieve superior performances in terms of OA, mAcc, and CD, comparing to SOTA methods that are specifically trained on ScanObjectNN with complete shapes. Notably, all of our jointly learned HSD models already outperform the most-recent hierarchy-based methods Point-PN and PointConT. This highlights the effectiveness and importance of the proposed method on the real-world setup. More results on ScanObjectNN without background are given in Table. V. Here, we primarily compare our methods to our baselines, except for the upstream’s reconstruction error without joint learning. It can be again observed that our proposed methods achieve lower CD than merely using a completion network SnowflakeNet.
| Model | OA(%) | mAcc(%) | CD |
| PointMAE [14] | 77.6 | 74.7 | 17.79 |
| PointMLP [12] | 77.7 | 75.1 | |
| Point-PN [36] | 79.0 | 76.5 | |
| PointConT [37] | 81.2 | 78.1 | |
| Our (PointNet++) | 79.1 | 76.1 | 17.95 |
| Our (CurveNet) | 80.1 | 77.0 | 17.76 |
| Our (PointMLP) | 81.6 | 78.9 | 17.82 |
| Our (PointNet++-HSD) | 81.3 | 79.0 | 16.79 |
| Our (CurveNet-HSD) | 81.6 | 78.8 | 16.74 |
| Our (PointMLP-HSD) | 82.7 | 80.8 | 16.35 |
| Model | OA(%) | mAcc(%) | CD | OA(%) | mAcc(%) | CD |
| SnowflakeNet [21] | - | - | 12.00 | - | - | 11.05 |
| Our (PointNet++) | 82.2 | 80.1 | 12.11 | 82.2 | 79.8 | 11.12 |
| Our (CurveNet) | 82.8 | 80.5 | 12.03 | 83.1 | 80.9 | 11.10 |
| Our (PointMLP) | 84.5 | 82.8 | 12.11 | 84.2 | 82.5 | 11.16 |
| Our (PointNet++-HSD) | 84.5 | 82.7 | 11.16 | 84.9 | 83.5 | 9.98 |
| Our (CurveNet-HSD) | 83.7 | 81.3 | 11.97 | 83.6 | 80.9 | 10.04 |
| Our (PointMLP-HSD) | 85.2 | 83.3 | 11.49 | 85.1 | 83.3 | 10.17 |
IV-D Quantitative Analysis
| Our | Level 1 | Level 2 | Level 3 | |
| (PointNet++) | OA/mAcc | OA/mAcc | OA/mAcc | CD |
| Baseline | 82.2/80.1 | - | - | 12.11 |
| DSN | 82.4/80.0 | 83.5/81.7 | 84.1/82.8 | 11.51 |
| SCL (HSD) | 82.8/80.5 | 83.6/81.7 | 83.9/83.0 | 11.25 |
| Full (HSD) | 82.8/81.0 | 83.9/82.0 | 84.5/82.7 | 11.17 |
Unlike images, one can only define probability distribution for continuously distributed variables, such as point clouds. In this section, we only measure information by discretization and relative information content in terms of KL divergence and cross entropy. See the supplemental for details of discretization. First, we visualize the information plane in Fig. 8(a), where both and gradually increase and eventually hit the maximum. This makes sense, since our MAE’s objective is to fully reconstruct from latent representations . Similar phenomenon also occurs in an autoencoder-based classifier [45]. This is also in line with other findings in [46, 47], where compression is not always necessary for good generalization performance. Second, since Fig. 8(a) merely differentiate the information of 3 levels separately, we quantify in terms of KL loss in Fig. 8(b). The differences ( and ) between companions and the teacher are decreasing, indicating that the information from the deepest level is flowing to previous levels, providing a stronger regularization. Finally, we utilize cross entropy to approximate in Fig. 8(c) , which suggests that the teacher shares most mutual information to real labels.
Specifically, we use 6 bins to digitize the information (i.e., network weights) from the 3 levels for the initial 100 epochs. Thereafter, with the probabilities of the given data and labels , mutual information and is therefore estimated via:
The procedure for calculating the mutual information and plotting the information plane is included in the source code.
IV-E Ablation Study
| Upstream | OA | mAcc | CD |
| PMP-Net++ [22] | - | - | 10.89 |
| Our (SnowflakeNet) | 84.5 | 82.7 | 11.16 |
| Our (PMP-Net++) | 85.1 | 83.5 | 9.73 |
| Model | mIOU(%) | cIOU(%) | CD |
| PointMLP [12] | 83.2 | 79.0 | 24.46 |
| PointNeXt [48] | 83.5 | 78.9 | |
| Our (PointNet++-HSD) | 83.3 | 79.7 | 24.18 |
| Our (PointMLP-HSD) | 83.5 | 80.2 | 24.14 |
| Model | Airplane | Bag | Cap | Car | Chair | Earph | Guitar | Knife | Lamp | Lapt | Motor | Mug | Pistol | Rocket | Skate | Table |
| PointMLP [12] | 79.3 | 78.4 | 78.7 | 72.8 | 87.9 | 67.1 | 89.8 | 86.4 | 79.9 | 95.1 | 60.6 | 91.9 | 79.4 | 60.6 | 73.7 | 82.7 |
| PointNeXt [48] | 80.1 | 74.5 | 80.2 | 72.6 | 88.1 | 72.6 | 89.7 | 86.4 | 82.8 | 94.6 | 55.0 | 91.3 | 79.7 | 59.6 | 72.3 | 82.1 |
| Our (PointNet++-HSD) | 80.1 | 80.6 | 78.8 | 72.3 | 87.8 | 74.5 | 91.1 | 85.9 | 79.8 | 95.0 | 58.1 | 92.8 | 81.4 | 60.6 | 72.9 | 83.2 |
| Our (PointMLP++-HSD) | 79.6 | 77.2 | 82.1 | 73.8 | 88.0 | 67.5 | 91.0 | 86.4 | 79.3 | 94.5 | 61.7 | 92.6 | 81.1 | 61.2 | 77.3 | 83.1 |
Sections IV-A and IV-C present the comparisons between our baselines and the models based on HSD, indicating the superior performance of the proposed HSD mechanism. The results also showcase the extendibility of the HSD mechanism in scenarios involving flawed or rapidly scanned point clouds. To validate the effectiveness of HSD, we conduct ablation studies with or without the self-distillation loss, on the PB_T50_RS variant (without background) of ScanObjectNN. Here, PointNet++ serves as the backbone for the downstream model, and the results are presented in Table VI. First, the performances of baseline and DSN (i.e., w/o. HSD as discussed in section III-D) are reported. In addition, the significance of the teacher’s code length is demonstrated by employing identical dimensions in the hyperspaces for the last two classification heads, as indicated by the same code length (SCL). Lastly, the full setting with HSD further regularizes for the optimization. These two observations suggest that a longer code length provides more information (i.e., ), and maximizing the mutual information provides stronger regularization.
Worth mentioning, the selection of MAE in our method is also plug-and-play, as shown in Table VII, where we use PMP-Net++ [22] as the upstream. Note that PMP-Net++ is simplified analogue to SnowflakeNet. With the same downstream network, i.e., PointNet++-HSD, Our model with a more powerful upstream again improves all three metrics. This verifies again HSD helps MAE to rearrange high-level details that cannot be reconstructed by MAE alone.
IV-F Part Segmentation
To demonstrate the generality of our framework, we compare our methods to the SOTAs of part segmentation in Table VIII. Since PointMLP and PointNeXt use both coordinate and normal information for training, we retrain them with only the coordinates (2,048 points) on the training set of ShapeNetPart, and test the result generated from the upstream. Both training and testing sets are normalized for consistency. Under an unfair condition, i.e., with merely one half of input (1,024 points) for training, our simplest model (PointNet++-HSD) already outperforms PointMLP. This suggests that our framework possesses the potential to elevate the performance of a model to that of a subsequent generation, and reaffirms the generalization ability of our framework for scattered point clouds on the part segmentation task. We elaborate the results of per-class segmentation IOU in Table IX Specifically, the per-class IOUs are obtained from the checkpoints of the best cIOU for fair comparisons.
Implementation of Part Segmentation. The loss computation and evaluation of incomplete point cloud segmentation are non-trivial. The per-point accuracy and mean IOU’s computation necessitate precise point order as in the ground truth data. Given that the generated point cloud lacks this precise order, direct computation of these metrics becomes unfeasible. However, the reconstruction loss presents an opportunity to roughly compute them, as it formulates the bidirectional mapping, albeit not on a per-point basis, between the predicted and ground truth point clouds. Based upon this mapping, the point index orders are aligned between the two point sets.
IV-G Failure Case
ModelNet40 encompasses a substantial number of classes that exhibit similarities in shape, e.g., table and bench, known as similar-class issue. We consider the similar-class issue as a partial failure case, especially when the sparsity level is low or medium. Another potential failure case is for extremely complex shapes, which may not be appropriately recovered.
V Conclusion
Recent state-of-the-art models have shown less effectiveness on incomplete point clouds. In this paper, we have proposed a cascaded network that tailors the upstream and downstream networks for analyses on scattered point clouds. In our method, MAE serves as a rudimentary form of object completion focusing on low-level features, while HSD further refines the understanding of this preliminary reconstruction focusing on high-level features. Our experiments demonstrate that our joint learning design performs good generalization on scattered point clouds, indicating that the reconstructed objects are beneficial for better environmental understanding, particularly in scenarios where perfect scanning is impractical during data collection. Moreover, our proposed hierarchical self-distillation improves the capacity of hierarchy-based classification models by maximizing the mutual information among layers, indicating that the optimization process is further regularized by the information flow.